# Gibbs polystability of Fano manifolds, stability thresholds and symmetry breaking

### v2

@@ 6. The $\mathcal{K}$–reduced probability measures on $X^{N}$
We continue with the setup introduced in Section 3. We thus fix a compact subgroup $\mathcal{K}$ of $\mathcal{G}$ and, given a log Fano manifold $(X,\Delta),$ we fix a $\mathcal{K}$–invariant positively curved metric on $K_{(X,\Delta)}.$ Recall that $\boldsymbol{m}_{N}$ denotes the corresponding moment map for the diagonal action of $\mathcal{K}$ on $X^{N}$, defined $X^{N}$ with values in the dual $\mathfrak{k}^{*}$ of the Lie algebra $\mathfrak{k}$ of $\mathcal{K}$ (Section 3.2.8).
@@
Before introducing the probabilistic setup we start with some preparations. First assume that 0 is a regular value of the moment map $\boldsymbol{m}_{N}.$ Given a measure $\nu$ on $X^{N}$ with a smooth density and a basis $e_{1},\dots,e_{r}$ of $\mathfrak{k}$, we will denote by $\nu_{\boldsymbol{m}_{N}}$ the corresponding Gelfand–Leray form on the submanifold $\left\{\boldsymbol{m}_{N}=0\right\}$ of $X^{N},$ uniquely determined by the following property:
@@ 6.1. Probabilistic and statistical mechanical setup
6.1.1. The Gelfand–Leray measure $\nu_{\boldsymbol{m}_{N}}^{(N)}$ and the $\mathcal{K}$–reduced canonical probability measure $\mu_{\boldsymbol{m}_{N}}^{(N)}$ and partition function $Z_{\boldsymbol{m}_{N}}.$
Given a log Fano manifold $(X,\Delta),$ denote by $\nu^{(N)}$ the measure on $X^{N}$ defined by
where $\det S^{(N)}$ denotes the section of $kK_{(X^{N},\Delta_{N})}\rightarrow X^{N}$ obtained by taking $s_{1},\dots,s_{N}$ in formula 1.3 to be a basis of $H^{0}(X,-kK_{(X,\Delta)})$ which is orthonormal with respect to the Hermitian product on $H^{0}(X,-kK_{(X,\Delta)})$ induced by the metric $\phi_{0}$ on $K_{(X,\Delta)}$ and the measure $\mu_{\phi_{0}}$ on $X.$ When 0 is a regular value of $\boldsymbol{m}_{N}$ we will denote by $\nu_{\boldsymbol{m}_{N}}^{(N)}$ the corresponding Gelfand–Leray measures, supported on $X^{N}.$ While this measure depends on the choice of basis in $\mathfrak{k},$ changing the basis only has the effect of multiplying $\nu_{\boldsymbol{m}_{N}}$ with a positive constant $C$ (the determinant of the corresponding change of bases matrix). Since $C$ is independent of $N$ this does not effect the large-$N$ limits in Conjecture 1.6 and Theorem 1.7.
When 0 is not a regular value of $\boldsymbol{m}_{N}$ we propose, tentatively, to define $\nu_{\boldsymbol{m}_{N}}$ as
@@
[…]
@@
When 0 is a regular value of $\boldsymbol{m}_{N}$ the previous definition coincides with the limit of the rhs of formula 6.2 as $\epsilon_{N}\rightarrow 0$ (for $N$ fixed). It would be interesting to know if, in general, this limit exists and if it coincides with the measure supported on the regular locus, defined by the Gelfand–Leray type measure on the regular locus.
Set
(6.3) $\mu_{\boldsymbol{m}_{N}}^{(N)}\coloneqq\nu_{\boldsymbol{m}_{N}}^{(N)}/Z_{\boldsymbol{m}_{N}},\,\,\,Z_{\boldsymbol{m}_{N}}\coloneqq\int\nu_{\boldsymbol{m}_{N}}^{(N)}.$
Assuming that $Z_{\boldsymbol{m}_{N}}<\infty$ the measure $\mu_{\boldsymbol{m}_{N}}^{(N)}$ defines a probability measure on $X^{N},$ that we shall call the $\mathcal{K}$–reduced canonical probability measure and the corresponding normalizing constant $Z_{\boldsymbol{m}_{N}}$ will be called the $\mathcal{K}$–reduced partition function.
@@ 6.1.2. The contracted setup
One could also define an alternative probability measures $(\mu^{(N)})^{\mathrm{c}}$ by replacing the Gelfand–Leray measure $\nu_{\boldsymbol{m}_{N}}^{(N)}$ with the contracted measure $(\nu^{(N)})^{\mathrm{c}}$ and the $\mathcal{K}$–reduced partition function $Z_{\boldsymbol{m}_{N}}$ with the contracted partition function, defined by
@@ 6.1.3. The $\mathcal{K}$–reduced Gibbs measures at inverse temperature $\beta$
$\nu_{\beta}^{(N)}\coloneqq e^{-\beta NE^{(N)}}\mu_{0}^{\otimes N},\,\,\,\,\,E^{(N)}(x_{1},\dots,x_{N})\coloneqq-(kN)^{-1}\log\left(\left\|\det S^{(N)}(x_{1},\dots,x_{N})\right\|_{\phi_{0}}^{2}\right).$
When $\beta=-1$ it coincides with the measure $\nu^{(N)}$ on $X^{N}$ (but $\nu_{\beta}^{(N)}$ depends on the fixed metric $\phi_{0}$ when $\beta\neq-1).$ The measures $\nu_{\beta}^{(N)}$ on $X^{N}$ induce a family of Gelfand–Leray measures $\left(\nu_{\beta}^{(N)}\right)_{\boldsymbol{m}_{N}}.$ Assuming that the $\mathcal{K}$–reduced partition function $Z_{\boldsymbol{m}_{N}}(\beta)$ at inverse temperature $\beta$ is finite,
$Z_{\boldsymbol{m}_{N}}(\beta)\coloneqq\int\left(\nu_{\beta}^{(N)}\right)_{\boldsymbol{m}_{N}}<\infty,$
@@
(6.5) $\mu_{\boldsymbol{m}_{N},\beta}^{(N)}\coloneqq\frac{\left(\nu_{\beta}^{(N)}\right)_{\boldsymbol{m}_{N}}}{Z_{\boldsymbol{m}_{N}}(\beta)}=\frac{e^{-\beta NE^{(N)}}\left(\mu_{0}^{\otimes N}\right)_{\boldsymbol{m}_{N}}}{Z_{\boldsymbol{m}_{N}}(\beta)}.$
When $\beta=-1$ this measure coincides with the $\mathcal{K}$–reduced canonical probability measure $\mu_{\boldsymbol{m}_{N}}^{(N)}$, defined in the previous section, and when $\beta=0$ it coincides with the classical microcanonical measure associated to the Hamiltonians $h_{1},\dots,h_{r},$ that is ubiquitous in the statistical mechanics literature.
@@
From a statistical mechanics perspective, the function $E^{(N)}$ plays the role of the $N$-particle energy (or Hamiltonian) and the corresponding Gibbs measure $\mu_{\boldsymbol{m}_{N},\beta}^{(N)}$ describes the corresponding equilibrium state of $N$ interacting particles confined to $\left\{\boldsymbol{m}_{N}=0\right\},$ at inverse temperature $\beta.$ The corresponding probability space $(X^{N},\mu_{\boldsymbol{m}_{N},\beta}^{(N)})$ may be viewed as a singular incarnation of the notion of mixed ensemble introduced in [54]. In particular, when $X=\mathbb{P}^{1},$ identified with the two-sphere and $\omega_{0}$ is the standard $\mathrm{SO}(3)$–invariant symplectic form on $X,$ $E^{(N)}$ coincides with the energy particle in Onsager’s statistical mechanical description of the 2D incompressible Euler equations, adapted to the two-sphere [67, 68]. The corresponding Gibbs measures $\mu_{\boldsymbol{m}_{N},\beta}^{(N)}$ are invariant under the Hamiltonian flow of $E^{(N)}$ on $X^{N}$ and, building on Theorem 1.7, we show in the companion paper [4] that the empirical measure $\delta_{N},$ viewed as a random variable on the probability space $(X^{N},\mu_{\boldsymbol{m}_{N},\beta}^{(N)})$ converges in law towards the $\mathrm{SO}(3)$–invariant probability measure on the two-sphere, for any $\beta>-1.$
@@
For any $\beta\in\mathbb{R}$ the probability measure $\mu_{\boldsymbol{m}_{N},\beta}^{(N)}$ is $\mathcal{K}$–invariant, when it is well-defined.
@@
First observe that for any $g\in\mathcal{G},$ $g^{*}\det S^{(N)}=\chi_{N}(g)\det S^{(N)}$ for a homomorphism $\chi_{N}\colon\mathcal{G}\rightarrow\mathbb{C}^{*}$ (see Lemma 6.9 below). It follows that the restriction of $\chi_{N}$ to $\mathcal{K}$ takes values in $S^{1}\subset\mathbb{C}^{*}.$ Indeed, since $\mathcal{K}$ is compact the image $\chi_{N}(\mathcal{G})$ is a compact subgroup of $\mathbb{C}^{*}$ and thus contained in the unique maximally compact subgroup $S^{1}$ of $\mathbb{C}^{*}.$ Since $\phi_{0}$ is $\mathcal{K}$–invariant, it follows that $\left\|\det S^{(N)}(x_{1},\dots,x_{N})\right\|_{\phi_{0}}^{2}$ is $\mathcal{K}$–invariant on $X^{N}.$ Finally, since the measure $\mu_{0}$ and $|\boldsymbol{m}|$ are also $\mathcal{K}$–invariant we deduce that $\mu_{\boldsymbol{m}_{N},\beta}^{(N)}$ is $\mathcal{K}$–invariant. ∎

### v3

@@ 6. The $\mathcal{K}$–reduced probability measures on $X^{N}$
We continue with the setup introduced in Section 3. Given a log Fano manifold $(X,\Delta),$we thus fix a compact subgroup $\mathcal{K}$ of $\mathcal{G}$ and a $\mathcal{K}$–invariant positively curved metric on $K_{(X,\Delta)}.$ Recall that $\boldsymbol{m}_{N}$ denotes the corresponding moment map for the diagonal action of $\mathcal{K}$ on $X^{N}$, taking values in the dual $\mathfrak{k}^{*}$ of the Lie algebra $\mathfrak{k}$ of $\mathcal{K}$ (Section 3.2.9).
Given a positive rational number $k$ such that $kK_{(X,\Delta)}$ defines a line bundle we denote by $N$ the dimension of $H^{0}(X,-kK_{(X,\Delta)})$ and by $\nu^{(N)}$ the measure on $X^{N}$ defined by
where $\det S^{(N)}$ denotes the section of $kK_{(X^{N},\Delta_{N})}\rightarrow X^{N}$ obtained by taking $s_{1},\dots,s_{N}$ in formula 1.3 to be a basis of $H^{0}(X,-kK_{(X,\Delta)})$ which is orthonormal with respect to the Hermitian product on $H^{0}(X,-kK_{(X,\Delta)})$ induced by the metric $\phi_{0}$ on $K_{(X,\Delta)}$ and the measure $\mu_{\phi_{0}}$ on $X.$
Lemma 6.1.
Given a log Fano manifold $(X,\Delta),$ there exists a real character $\chi_{N}$ on $\mathcal{G}$ such that $g^{*}\nu_{N}=\chi_{N}(g)\nu_{N}$ for any $g\in\mathcal{G}.$ More precisely, given $W$ in the Lie algebra of $\mathcal{G}$,
(6.1) $\chi_{N}(e^{tW})=e^{tN2\text{Re}\left(\text{Fut}_{(X,\Delta),k}(W)\right)},$
for $t\in\mathbb{C},$ where $\text{Fut}_{(X,\Delta),k}(W)$ denotes the quantized Futaki invariant at level $k.$ Moreover, if $k$ is a positive integer and $k\Delta$ is a divisor, then $\text{Fut}_{(X,\Delta),k}=0$ iff $\det S^{(N)}$ is $\mathcal{G}$-invariant.
Proof.
First assume that $k$ is a positive integer and $k\Delta$ is a divisor. Then $\mathcal{G}$ acts naturally on $(X,-kK_{(X,\Delta)})$ and, as a consequence, also on $H^{0}(X,-kK_{(X,\Delta)}).$ By basic linear algebra, $g^{*}\nu_{N}=\left|\det A(g)\right|^{-2/k}\nu_{N}$ where $A(\cdot)$ denotes the representation of $\mathcal{G}$ on $H^{0}(X,-kK_{(X,\Delta)})$ and $\det A(g)$ defines a $\mathbb{C}^{*}$-valued algebraic character on $\mathcal{G}.$ Hence, $\chi_{N}(g)\coloneqq\left|\det A(g)\right|^{-2/k}$ is a real-valued character on $\mathcal{G}$ and $\chi_{N}(e^{tW})=e^{tN2\text{Re}\left(\text{Fut}_{(X,\Delta),k}(W)\right)}$ by the very definition of $\text{Fut}_{(X,\Delta),k}(W).$ In particular, if $\text{Fut}_{(X,\Delta),k}=0$ then $\det A(g)$ is a holomorphic function on $\mathcal{G}$ taking values in $S^{1}.$ But thus implies that $\det A(g)$ is identically equal to $1,$ showing that $\det S^{(N)}$ is is $\mathcal{G}$-invariant. Finally, consider the case when it is only assumed that $kK_{(X,\Delta)}$ defines a line bundle. Then $\mathcal{G}$ still acts naturally on the real line bundle over $X^{N}$ defined as the the top exterior power of the real cotangent bundle of $X^{N},$ having $\left|\det S^{(N)}\right|^{-2/k}$ as a smooth section. We can thus proceed essentially as before. ∎
Define the energy per NN particles as the following lsc function on $X^{N}$
$E^{(N)}(x_{1},\dots,x_{N})\coloneqq-(kN)^{-1}\log\left(\left\|\det S^{(N)}(x_{1},\dots,x_{N})\right\|_{\phi_{0}}^{2}\right)$
By 2.1,
(6.2) $E^{(N)}(x_{1},\dots,x_{N})=\infty\iff\exists s\in H^{0}(X,-kK_{(X,\Delta)})-\{0\}\text{ such that }s(x_{1})=\cdots=s(x_{N})$
As shown in [13], as $N\rightarrow\infty,$ $E^{(N)}$ converges towards the pluricomplex energy $E$ on $\mathcal{P}(X),$ in the sense of Gamma-convergence. The first part of the following result can thus be viewed as an analog of Prop 3.4 for finite $N$ and the second part is an analog of the fact that the unique minimizer $\omega_{0}^{n}/V$ of the functional $E$ on $\mathcal{P}(X)$ is in $\mathcal{P}(X)_{0},$ if the Futaki character of $(X,\Delta)$ vanishes (by Lemma 3.3).
Proposition 6.2.
Let $(X,\Delta)$ be a log Fano manifold. Assume that the quantized Futaki character of $(X,\Delta)$ at level $k$ vanishes. Then, under the diagonal action of $\mathcal{G}$ of $X^{N},$
$E^{(N)}(x_{1},\dots,x_{N})<\infty\implies(x_{1},...,x_{N})\,\text{is semi-stable }$
As a consequence, the closure of the $\mathcal{G}$-orbit of $(x_{1},...,x_{N})$ intersects $\left\{\boldsymbol{m}_{N}=0\right\}.$ Morover, any configuration $(x_{1},\dots,x_{N})_{*}$ minimizing $E^{(N)}$ satisfies $\boldsymbol{m}_{N}((x_{1},\dots,x_{N})_{*})=0.$ Additionally, if $H^{0}(X,-kK_{(X,\Delta)}\otimes T^{*}X)\neq\{0\}$, then $E^{(N)}(x_{1},\dots,x_{N})<\infty$ implies that $(x_{1},...,x_{N})$ has a finite stabilizer. In particular, then $\boldsymbol{m}_{N}$ is a submersion at $(x_{1},\dots,x_{N})_{*}.$
Proof.
First assume that $k$ is a positive integer such that $k\Delta$ is a divisor. Then $\mathcal{G}$ naturally acts on $kK_{(X,\Delta)}$ and the section $\det S^{(N)}$ is $G$-invariant, by Lemma 6.1. Hence, if $\det S^{(N)}(x_{1},\dots,x_{N})\neq 0,$ then $(x_{1},\dots,x_{N})$ is semi-stable, by the very definition of semi-stability. Next, if $\boldsymbol{x}:=(x_{1},\dots,x_{N})_{*}$ minimizes $E^{(N)}$ then it follows, in particular, that $\boldsymbol{x}$ minimizes the restriction of $E^{(N)}$ to the orbit $\mathcal{G}\boldsymbol{x}.$ Note that since $E^{(N)}$ is $G$-invariant its restriction $f$ to $\mathcal{G}\boldsymbol{x}$ defines a Kempf-Ness, i.e. up to scaling,
$f(g\boldsymbol{x})=-\log\left\|gS_{\boldsymbol{x}}\right\|(g\boldsymbol{x})$
for a non-zero vector $S_{\boldsymbol{x}}$ in the complex line over $\boldsymbol{x},$ which in this case is $(-kK_{X^{N_{k}}})_{|\boldsymbol{x}}.$ But in general a point $\boldsymbol{x}$ is a minimum of the Kempf-Ness function on $\mathcal{G}\boldsymbol{x}$ iff the corresponding moment map vanishes at $\boldsymbol{x}$ [83]. In the general case when it is only assumed that $kK_{(X,\Delta)}$ is defined as a line bundle, one can replace $\det S^{(N)}$ with a tensor power $(\det S^{(N)})^{\otimes m}$ for $m$ a sufficently divisible positive integer and proceed essentially as before. Finally, assume that $H^{0}(X,-kK_{(X,\Delta)}\otimes T^{*}X)\neq\{0\}.$ Assume that $E^{(N)}(x_{1},\dots,x_{N})\neq\infty$ and - in order to get a contradiction - that the stabilizer of $(x_{1},\dots,x_{N})$ is infinite. Then there exists a non-trivial holomorphic vector field $V$ on $X$ vanishing at all points $x_{1},...,x_{N}.$ Take an element $S_{k}$ in $H^{0}(X,-kK_{(X,\Delta)}\otimes T^{*}X)-\{0\}.$ Contracting $S_{k}$ with $V$ then gives a non-trivial holomorphic section of $kK_{(X,\Delta)}$ vanishing at all points $x_{1},...,x_{N},$ which contradicts $E^{(N)}(x_{1},\dots,x_{N})\neq\infty.$ Finally, we recall that, in general, the moment map is a submersion at $\boldsymbol{x}$ iff the stabilizier is finite iff $\boldsymbol{x}$ is stable. ∎
Note that the assumption $H^{0}(X,-kK_{(X,\Delta)}\otimes T^{*}X)\neq\{0\}$ is satisfied when $k$ is sufficently large, since $K_{(X,\Delta)}$ is ample. In general, this assumption is needed to get finite stabilizers, as illustrated in the following example.
Example 6.3.
Let $(X,\Delta)=(\mathbb{P}^{1},0).$ In this case $k\in\mathbb{Z}_{+}/2,$ $N=2k+1=2,3,$... and $H^{0}(X,-kK_{(X,\Delta)}\otimes T^{*}X)\neq\{0\}$ holds iff $k\geq 1,$ i.e. $N\geq 3.$ For any $N\geq 2$, $E^{(N)}(x_{1},\dots,x_{N})\neq 0$ iff all $x_{i}$ are distinct. Hence, $\left\{E^{(N)}(x_{1},\dots,x_{N})\neq\infty\right\}$ is a subset of $(X^{N})_{ss}$ , which is a proper subset when $N>3$ (by the condition 2.10). Consider now the case when $k=1/2$ (i.e. $N=2)$ and take $\phi_{0}$ to be the Fubini-Study metric i.e. $\phi_{0}$ is an $SO(3)$-invariant metric. Identifying $\mathbb{P}^{1}$ with the unit-sphere in $\mathbb{R}^{3},$ $\left\|\det S^{(2)}(x_{1},x_{2})\right\|_{\phi_{0}}$ is proportional to $|x_{1}-x_{2}|$ and $\boldsymbol{m}_{2}=(x_{1}+x_{2})/2.$ Hence, $(x_{1},x_{2})$ minimizes $E^{(2)}(x_{1},x_{2})=0$ iff $x_{1}$and $x_{2}$ are anti-podal points, which is equivalent to $\boldsymbol{m}_{2}(x_{1},x_{2})=0.$ This is also a consequence of the previous proposition. However, in this case the stabilizer is infinite (and, as consequence, $\boldsymbol{m}_{2}$ is not a submersion at $(x_{1},x_{2}).$). Indeed, any two distinct points on $\mathbb{P}^{1}$ are fixed by $\mathbb{C}^{*}-a$ction. Note also that $N=1$ on $\mathbb{P}^{1}$ is not covered by the previous proposition, which is consistent with the fact that there are no semi-stable points in this case (since the condition 2.10 is never satisfied). In particular, the moment map has no zeroes.
6.1. Gelfand–Leray measures and contracted measures
@@
Before introducing the probabilistic setup we start with some general definitions. Assume that 0 is a regular value of the moment map $\boldsymbol{m}_{N}.$ Given a measure $\nu$ on $X^{N}$ with a smooth density and a basis $e_{1},\dots,e_{r}$ of $\mathfrak{k}$, we will denote by $\nu_{\boldsymbol{m}_{N}}$ the corresponding Gelfand–Leray form on the submanifold $\left\{\boldsymbol{m}_{N}=0\right\}$ of $X^{N},$ uniquely determined by the following property:
@@ 6.2. Probabilistic and statistical mechanical setup
6.2.1. The Gelfand–Leray measure $\nu_{0}^{(N)}$ and the $\mathcal{K}$–reduced canonical probability measure $\mu_{0}^{(N)}$ and partition function $Z_{N,0}.$
Assume that 0 is a regular value of $\boldsymbol{m}_{N}.$ We then denote by $\nu_{0}^{(N)}$ the Gelfand–Leray measure supported on $\left\{\boldsymbol{m}_{N}=0\right\}\subset X^{N},$ corresponding to the measure $\nu^{(N)}$ on $X^{N}.$ While this measure depends on the choice of basis in $\mathfrak{k},$ changing the basis only has the effect of multiplying $\nu_{0}^{(N)}$ with a positive constant $C$ (the determinant of the corresponding change of bases matrix). Since $C$ is independent of $N$ this does not effect the large-$N$ limits in Conjecture 1.8 and Theorem 1.9. Set
(6.4) $\mu_{0}^{(N)}\coloneqq\nu_{0}^{(N)}/Z_{N,0},\,\,\,Z_{N,0}\coloneqq\int\nu_{0}^{(N)}.$
Assuming that $Z_{N,0}<\infty$ the measure $\mu_{0}^{(N)}$ defines a probability measure on $X^{N},$ that we shall call the $\mathcal{K}$–reduced canonical probability measure and the corresponding normalizing constant $Z_{N,0}$ will be called the $\mathcal{K}$–reduced partition function.
6.2.2. The thickened setup
When 0 is not a regular value of $\boldsymbol{m}_{N}$ we replace the measure $\nu_{0}^{(N)}$ with
@@ 6.2.2. The thickened setup
for an appropriate sequence of positive numbers $\epsilon_{N},$ called the thickening sequence. The corresponding normalization constants $Z_{N,\epsilon_{N}}$ and probability measurse $\mu_{\epsilon_{N}}^{(N)}$ will be called the thickened partition functions and probability measures.
@@
When 0 is a regular value of $\boldsymbol{m}_{N}$, $\nu_{0}^{(N)}$ coincides with the limit of $\nu_{\epsilon_{N}}^{(N)}$ as $\epsilon_{N}\rightarrow 0$ (for $N$ fixed). It would be interesting to know if, when $\{\boldsymbol{m}_{N}=0\}$ is non-empty, this limit exists and if it coincides with the measure supported on the regular locus of $\{\boldsymbol{m}_{N}=0\},$ defined by the Gelfand–Leray type measure on the regular locus.
@@ 6.2.3. The contracted setup
One could also define an alternative probability measures $(\mu^{(N)})^{\mathrm{c}}$ by replacing the Gelfand–Leray measure $\nu_{0}^{(N)}$ with the contracted measure $(\nu^{(N)})^{\mathrm{c}}$ and the $\mathcal{K}$–reduced partition function $Z_{N,0}$ with the contracted partition function, defined by
@@ 6.2.4. The $\mathcal{K}$–reduced Gibbs measures at inverse temperature $\beta$
$\nu_{\beta}^{(N)}\coloneqq e^{-\beta NE^{(N)}}\mu_{0}^{\otimes N}.$
When $\beta=-1$ it coincides with the measure $\nu^{(N)}$ on $X^{N}$ (but $\nu_{\beta}^{(N)}$ depends on the fixed metric $\phi_{0}$ when $\beta\neq-1).$ The measures $\nu_{\beta}^{(N)}$ on $X^{N}$ induce a family of Gelfand–Leray measures $\left(\nu_{\beta}^{(N)}\right)_{\boldsymbol{m}_{N}}.$ Assuming that the $\mathcal{K}$–reduced partition function $Z_{N,0}(\beta)$ at inverse temperature $\beta$ is finite,
$Z_{N,0}(\beta)\coloneqq\int\left(\nu_{\beta}^{(N)}\right)_{\boldsymbol{m}_{N}}<\infty,$
@@
(6.7) $\mu_{0,\beta}^{(N)}\coloneqq\frac{\left(\nu_{\beta}^{(N)}\right)_{\boldsymbol{m}_{N}}}{Z_{N,0}(\beta)}=\frac{e^{-\beta NE^{(N)}}\left(\mu_{0}^{\otimes N}\right)_{\boldsymbol{m}_{N}}}{Z_{N,0}(\beta)}.$
When $\beta=-1$ this measure coincides with the $\mathcal{K}$–reduced canonical probability measure $\mu_{0}^{(N)}$, defined in the previous section, and when $\beta=0$ it coincides with the classical microcanonical measure associated to the Hamiltonians $h_{1},\dots,h_{r},$ that is ubiquitous in the statistical mechanics literature (especially when $r=1)$. In general, $Z_{N,0}(\beta)$ is finite when $\beta>-\alpha(X,\Delta)$ (by the inequality 2.5). But if Conjecture 1.2 is correct, then this bound can be considerably improved.
@@
From a statistical mechanics perspective, the function $E^{(N)}$ plays the role of the $N$-particle energy (or Hamiltonian) and the corresponding Gibbs measure $\mu_{0,\beta}^{(N)}$ describes the corresponding equilibrium state of $N$ interacting particles confined to $\left\{\boldsymbol{m}_{N}=0\right\},$ at inverse temperature $\beta.$ The corresponding probability space $(X^{N},\mu_{0,\beta}^{(N)})$ may be viewed as a singular incarnation of the notion of mixed ensemble introduced in [70]. In particular, when $X=\mathbb{P}^{1},$ identified with the two-sphere and $\omega_{0}$ is the standard $\mathrm{SO}(3)$–invariant symplectic form on $X,$ $E^{(N)}$ coincides with the energy particle in Onsager’s statistical mechanical description of the 2D incompressible Euler equations, adapted to the two-sphere [87, 88]. The corresponding Gibbs measures $\mu_{0,\beta}^{(N)}$ are invariant under the Hamiltonian flow of $E^{(N)}$ on $X^{N}$ and, building on Theorem 1.9, we show in the companion paper [5] that the empirical measure $\delta_{N},$ viewed as a random variable on the probability space $(X^{N},\mu_{0,\beta}^{(N)})$ converges in law towards the $\mathrm{SO}(3)$–invariant probability measure on the two-sphere, for any $\beta>-1.$
@@
For any $\beta\in\mathbb{R}$ the probability measure $\mu_{0,\beta}^{(N)}$ is $\mathcal{K}$–invariant, when it is well-defined.
@@
First observe that for any $g\in\mathcal{G},$ $g^{*}\det S^{(N)}=\chi_{N}(g)\det S^{(N)}$ for a homomorphism $\chi_{N}\colon\mathcal{G}\rightarrow\mathbb{C}^{*}$ (see Lemma 6.1 below). It follows that the restriction of $\chi_{N}$ to $\mathcal{K}$ takes values in $S^{1}\subset\mathbb{C}^{*}.$ Indeed, since $\mathcal{K}$ is compact the image $\chi_{N}(\mathcal{G})$ is a compact subgroup of $\mathbb{C}^{*}$ and thus contained in the unique maximally compact subgroup $S^{1}$ of $\mathbb{C}^{*}.$ Since $\phi_{0}$ is $\mathcal{K}$–invariant, it follows that $\left\|\det S^{(N)}(x_{1},\dots,x_{N})\right\|_{\phi_{0}}^{2}$ is $\mathcal{K}$–invariant on $X^{N}.$ Finally, since the measure $\mu_{0}$ and $|\boldsymbol{m}|$ are also $\mathcal{K}$–invariant we deduce that $\mu_{0,\beta}^{(N)}$ is $\mathcal{K}$–invariant. ∎